OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. y = 10x - 9 y = x + 18 A. (4, 31) B. (3, 21) C. (1, 1) D. (2, 11)

Parth (parthkohli):

http://openstudy.com/users/parthkohli#/updates/4fa7aa68e4b029e9dc388281 See this tutorial, after that, see my method: \(\Large \color{purple}{\rightarrow x + 18 = 10x - 9 }\) \(\Large \color{purple}{\rightarrow x + 27 = 10x }\) \(\Large \color{purple}{\rightarrow 9x = 27 }\) \(\Large \color{purple}{\rightarrow x = 3 }\) Now, y = x + 18. We can infer: \(\Large \color{purple}{\rightarrow y = 3 + 18 }\) \(\Large \color{purple}{\rightarrow y = 21 }\) Hope this helps :)

OpenStudy (anonymous):

yeaa it do I just be wanting tosee if I have the correct answers